A solid copper cube has an edge length of 85.5 cm.
Answer #1

Bulk modulus of copper = B = 1.4 x 1011 N/m2

Initial edge length of the copper cube = L1 = 85.5 cm = 0.855 m

Final edge length of the copper cube = L2 = 85 cm = 0.85 m

Initial volume of the cube = V1

V1 = L13

V1 = (0.855)3

V1 = 0.625 m3

New volume of the cube = V2

V2 = L23

V2 = (0.85)3

V2 = 0.6141 m3

Change in volume of the cube = \DeltaV

\DeltaV = V1 - V2

\DeltaV = 0.625 - 0.6141

\DeltaV = 0.0109 m3

Stress applied to the cube = P

B = \frac{P}{\Delta V/V}

P = \frac{B\Delta V}{V}

(1.4 * 1011) (0.0109) (0.625)

P = 2.441 x 109 N/m2

Stress applied to the copper cube to reduce its edge length from 85.5 cm to 85 cm = 2.441 x 109 N/m2